OpenStudy (anonymous):

The additive inverse of a number is the number that, when added to the number, yields zero. True False

OpenStudy (anonymous):

True. additive inverse of a number is the number's negative

OpenStudy (anonymous):

let the number =a additive inverse of a = -a a + (-a) = a -a =0